# Demonstrating CavePI: Autonomous Exploration of Underwater Caves by Semantic Guidance

### v1

@@ Abstract
Enabling autonomous robots to safely and efficiently navigate, explore, and map underwater caves is of significant importance to water resource management, hydrogeology, archaeology, and marine robotics. In this work, we demonstrate the system design and algorithmic integration of a visual servoing framework for semantically guided autonomous underwater cave exploration. We present the hardware and edge-AI design considerations to deploy this framework on a novel AUV (Autonomous Underwater Vehicle) named CavePI. The guided navigation is driven by a computationally light yet robust deep visual perception module, delivering a rich semantic understanding of the environment. Subsequently, a robust control mechanism enables CavePI to track the semantic guides and navigate within complex cave structures. We evaluate the system through field experiments in natural underwater caves and spring-water sites and further validate its ROS (Robot Operating System)-based digital twin in a simulation environment. Our results highlight how these integrated design choices facilitate reliable navigation under feature-deprived, GPS-denied, and low-visibility conditions.
@@ 1 Introduction
Autonomous underwater cave exploration is an important area of research, providing valuable insights into some of the planet’s least explored ecosystems to understand and manage marine resources. Underwater caves, in particular, hold immense potential for advancing our knowledge of archaeology, hydrology, and geology [23]. The formations and sediments within these caves serve as critical records of historical climate and geology events while playing a key role in monitoring groundwater flow within Karst topographies, which supply freshwater to nearly a quarter of the global population [14]. However, these environments are difficult to access and often hazardous for human divers due to their confined spaces, complex geometries, and lack of natural light [36, 5]. To this end, AUVs present a promising solution, offering safe and efficient exploration while minimizing risks and improving precision [1].
@@
However, fully autonomous exploration inside underwater caves presents significant challenges, navigation is difficult without GPS, while poor visibility, backscattering, and silt disturbances hinder perception. Thus, uninformed exploration, as often done by AUVs in open-water ecological surveying and mapping tasks [28], is quite challenging. For autonomous navigation inside underwater caves [42, 59], it is important to be aware of the entry and exit directions as well as a semantic understanding of existing navigational markers, when available. More specifically, underwater caves explored by humans are augmented by a line (string), referred to as the caveline [13], which extends from the entrance through all the major sections of the cave. Other, navigational markers such as arrows and cookies [1] indicate the closest path to the exit (arrows’ directions) and the presence of landmarks and/or divers (cookies). The caveline, together with the navigational markers, provides a 1D retraction of the 3D environment, indicating the orientation and distance from the entrance of the primary passages.
@@
In this paper, we demonstrate the system design and algorithmic integration of CavePI, a novel AUV specifically developed for navigating underwater caves using semantic guidance from cavelines and other navigational markers; see Fig. 1. The compact design and 4-degree-of-freedom (4-DOF) motion model enable safe traversal through narrow passages with minimal silt disturbance. Designed for single-person deployment, CavePI features a forward-facing camera for visual servoing and a downward-facing camera for caveline detection and tracking, effectively minimizing blind spots around the robot. A Ping sonar provides sparse range data to ensure robust obstacle avoidance and safe navigation within the caves. Additionally, two onboard lights enhance visibility in low-light conditions, complemented by visual filters for improved perception. The computational framework is powered by two single-board computers (SBCs): a Jetson Nano for perception, and a Raspberry Pi-5 for planning and control.
@@
For visual servoing, we integrate a caveline detector that performs real-time pixel-level segmentation of the caveline using the robot’s down-facing camera stream. While prior studies have proposed vision transformer (ViT)-based frameworks such as CL-ViT [59] and CaveSeg [1] for caveline detection, these approaches are computationally intensive and unsuitable for the resource-constrained SBCs. To address this, we adopt more efficient feature-extractor networks: MobileNetV3 [18] and ResNet101 [16] – combined with a lightweight DeeplabV3 [8] as the segmentation head. By leveraging GPU-accelerated computation on a Jetson Nano SBC, our system achieves a maximum inference rate of $18.2$ frames per second (FPS) while ensuring robust segmentation performance. The resulting segmentation map is post-processed to extract caveline contours, which are used to guide a PID-based Pure Pursuit controller for precise heading adjustments for smooth tracking-by-detection [45].
@@
Following extensive in-house testing and refinement, CavePI is deployed in diverse real-world environments, including $1.5$ m-$6$ m deep spring waters and $12$ m-$30$ m deep natural underwater grottos and caves. These outdoor settings pose unique perception and navigation challenges, such as strong currents and turbidity in open waters, as well as noisy, low-light conditions within overhead environments. These deployments validate the robustness of CavePI, demonstrating its capability to perform long-term autonomous missions in complex underwater environments.
@@ 2.1 Autonomous Underwater Cave Exploration
Exploring underwater caves remains a perilous and demanding endeavor, even for experienced scuba divers, due to confined spaces, little to no natural lights, and a high risk of disorientation [5]. Thus, ROVs and AUVs play a crucial role in tasks such as mapping and sample collection [46]. While SOTA AUVs are proficient in executing fully autonomous missions, their deployments have primarily been constrained to open water environments [43], where obstacles are minimal. Underwater caves, however, present unique and formidable navigation challenges, including overhead and surrounding obstacles, narrow passageways, and complex topologies. Early efforts to address these challenges include the work of Mallios et al. [27], where manually operated AUVs collected acoustic data for offline mapping. Similarly, Weidner et al. [56, 55] utilized stereo camera systems to generate high-resolution 3D reconstructions of cave walls, floors, and ceilings.
@@
Vision-based state estimation and navigation in underwater caves pose significant challenges due to factors such as lighting variability, backscattering effects, and image degradation [21]. To overcome these obstacles, Rahman et al. [38, 40] proposed a data fusion framework that integrates visual, acoustic, inertial, and water depth measurements, enabling robust trajectory estimation and sparse representations of underwater cave environments. Advances such as shadow-based mapping [39], contour extraction [29], and real-time stereo reconstruction [54] have further enhanced the density and fidelity of cave mapping techniques. More recently, visual learning-based approaches have shown promise for autonomous visual servoing within underwater caves [1, 59]. Other contemporary works with the Sunfish autonomous underwater vehicle (AUV) [42] has demonstrated effective navigation capabilities in these environments by employing Doppler Velocity Log (DVL)-based dead reckoning and sonar-based SLAM algorithms to navigate and map complex underwater cave systems.
@@ 2.2 Robot Navigation by Semantic Guidance
For navigating underwater caves safely, learning-based frameworks have been developed for detecting and following cavelines [59] and other navigation markers [1]. However, these computationally intensive models require further optimization to achieve real-time deployments [30]. Additionally, their onboard semantic segmentation performance and decision-making capabilities remain underexplored. We address this gap by conducting comprehensive real-world evaluations using our CavePI robotic platform, demonstrating its effectiveness for real-time semantic segmentation and navigation in underwater cave environments.
@@ 3.1 Sensor Bay: Acoustic-Optic Perception Subsystem
The CavePI platform includes visual and acoustic sensors: a front-facing fisheye camera, a downward-facing low-light camera, and a Ping2 active sonar. The fisheye camera, housed within a transparent dome at the head of the AUV, captures forward-facing visuals with a $160^{\circ}$ field-of-view (FOV) and outputs a video feed at $1920\times 1080$ resolution. It is worth noting that the cylindrical enclosure is a $6^{\prime\prime}$ tube while the dome has a $4^{\prime\prime}$ diameter. A custom interface is built to connect the two; see section 5.1 for more details. The low-light camera, mounted inside the computational enclosure, captures downward-facing visuals with an $80^{\circ}\times 64^{\circ}$ FOV, also at the same resolution. Additionally, a Ping2 sonar altimeter-echosounder from Blue Robotics™ is mounted on the underside of the robot; the sonar has a range of $100$ meters, a depth rating of $300$ meters, and a resolution of $0.5\%$ of the range, allowing it to detect obstacles in the surrounding environment beneath CavePI. These sensory components collectively provide robust environmental awareness for autonomous navigation in challenging underwater environments.
@@ 3.2 Computational Bay
The enclosure also contains the battery compartment, voltage regulators, electronic speed controllers (ESCs), and a Bar-30 pressure sensor. The battery compartment holds a $14.8$ V ($18$ Ah) battery pack, regulated to power internal components (e.g., cameras, computers) and external components (e.g., thrusters, sonar). Each thruster is controlled by an ESC, which drives the three-phase brushless motor using PWM signals from the Pixhawk. The Bar-30 sensor provides high-precision pressure readings with a resolution of $0.2$ mbar and an accuracy of $2$ mm, with a working depth of up to $300$ meters. This pressure data is processed to determine CavePI’s underwater depth, ensuring reliable and accurate interoceptive perception during operations.
@@ 3.3 Locomotion Bay: Middleware Integration
The CavePI autonomous underwater vehicle (AUV) is designed for low-power operation and integrates a portable ROS2 framework to support application-specific perception, planning, and control modules. As depicted in Fig. 4, the detector node acquires visual data from the two cameras to identify the caveline for navigation. In the absence of GPS underwater, the system employs SVIn (sonar-visual-inertial navigation) [41] packages to estimate the AUV’s position relative to the detected caveline. The mission planner node then integrates the caveline information with the estimated position data to generate subsequent waypoints for the mission. Finally, the autopilot controller node utilizes these waypoints, along with the detected caveline, positional data, and depth readings from the Bar-30 sensor, to generate precise actuation signals for the thrusters, enabling accurate movements and depth control. Additionally, CavePI can function as an ROV through an optional tether-based teleoperation module. This module transmits user commands from a joystick to the onboard Raspberry Pi-5, which processes the inputs and relays them to the thrusters for manual control.
@@ 3.4 CavePI Digital Twin
Beyond replicating caveline following experiments, we utilize the DT system for preliminary testing and fine-tuning of new control algorithms. It also enables the simulation of complex cave scenes, such as narrow passages, dead ends, and sharp turns. Conducting repeated real-world experiments in such scenarios to improve the control system can be logistically demanding where the simulation offers an efficient alternative for extensive evaluation and fine-tuning.
@@ 4 CavePI Navigation Pipeline
The CavePI AUV is designed to autonomously navigate underwater by following a caveline and other navigation markers. However, a caveline appears a few pixels wide in the bottom-facing camera’s FOV and is significantly challenging to detect in noisy low-light conditions. Moreover, the caveline is often obscured or blends into the background, requiring a fine-tuned visual learning pipeline for reliable operation.
@@ 4.1 Semantic Guidance: Tracking by Detection
The model selection process plays a crucial role in optimizing accuracy and efficiency on resource-constrained edge devices. While object detection and semantic segmentation are two established techniques in visual learning, they produce different interpretations of the target RoI. Object detection models generate bounding boxes around the target, which, given the caveline’s irregular shape and orientation, can encompass significant background regions, complicating the estimation of the heading angle. To address this, we opt for semantic segmentation, which provides pixel-level contours of the caveline as a more precise tracking-by-detection.
@@
Model architecture and design choices. Onboard resource constraints of CavePI significantly influence the choice of model architecture. The Jetson Nano, with its shared memory architecture, allocates GPU memory dynamically from a limited $2$ GB pool, which must also accommodate the dual camera feeds and Nano-Pi control signal communication. Considering these limitations, we initially selected the MobileNetV3 (Large) backbone [18] for feature extraction. This lightweight architecture feeds extracted features into a DeepLabV3 [8] head to generate binary segmentation maps. While the MobileNetV3-DeepLabV3 combination is computationally efficient, requiring only about $314$ MB GPU memory for $11$ million parameters, it proved insufficient for accurately segmenting the thin caveline in turbid water conditions. To address this limitation, we adopted a heavier backbone, ResNet101 [16], while retaining the DeepLabV3 head. This model achieved slightly better performance, as reported in Table 1, albeit at the cost of reduced inference speed. Both models were initialized with pre-trained weights and fine-tuned on a custom dataset to enhance caveline detection accuracy in challenging underwater environments.
Dataset and model training. For this study, we utilized the open-source CL-ViT dataset, which includes pixel-level annotations of cavelines [59]. It comprises $3,150$ RGB images collected from three distinct underwater cave systems. To enhance the dataset’s diversity and variance, we manually annotated an additional $150$ images from our laboratory testbed, resulting in a combined dataset of $3,300$ images. The dataset was divided into a training set of $3,200$ images and a validation set of $100$ images. Both models were fine-tuned on this dataset over a maximum of $50$ epochs using the following training configurations: the Adam optimizer with an initial learning rate of $0.001$, a cosine learning rate scheduler with $T_{max}=20$, and a cross-entropy loss function. This training strategy was designed to optimize these models’ performances in diverse underwater environments.
Benchmark comparison. The fine-tuned models are evaluated on the open-source CL-Challenge dataset [59] that contains caveline annotations for $200$ noisy low-light images from different cave systems. Using the lighter MobileNetV3 backbone, our model achieves a mean intersection-over-union (mIoU) of $48.95\%$, whereas the considerably heavier ResNet101 model demonstrates only a marginal improvement with a mIoU of $50.48\%$. For reference, the highest mIoU score reported on this dataset is $58.3\%$ [59].

### v3

@@ Abstract
Enabling autonomous robots to safely and efficiently navigate, explore, and map underwater caves is of significant importance to water resource management, hydrogeology, archaeology, and marine robotics. In this work, we demonstrate the system design and algorithmic integration of a visual servoing framework for semantically guided autonomous underwater cave exploration. We present the hardware and edge-AI design considerations to deploy this framework on a novel AUV (Autonomous Underwater Vehicle) named CavePI. The guided navigation is driven by a computationally light yet robust deep visual perception module, delivering a rich semantic understanding of the environment. Subsequently, a robust control mechanism enables CavePI to track the semantic guides and navigate within complex cave structures. We evaluate the CavePI system through field experiments in natural underwater caves and spring-water sites, and further validate its ROS (Robot Operating System)-based digital twin in a simulation environment. Our results highlight how these integrated design choices facilitate reliable navigation under feature-deprived, GPS-denied, and low-visibility conditions. The system design, code, and data are available on the project website: https://github.com/uf-robopi/CavePI_AUV/.
@@ 1 Introduction
Autonomous underwater cave exploration is an important area of research, providing valuable insights into some of the planet’s least explored ecosystems to understand and manage marine resources. Underwater caves, in particular, hold immense potential for advancing our knowledge of archaeology, hydrology, and geology [26]. The formations and sediments within these caves serve as critical records of historical climate and geology events while playing a key role in monitoring groundwater flow within Karst topographies, which supply freshwater to nearly a quarter of the global population [15]. However, these environments are difficult to access and often hazardous for human divers due to their confined spaces, complex geometries, and lack of natural light [40, 5]. To this end, Autonomous Underwater Vehicles (AUVs) present a promising solution, offering safe and efficient exploration while minimizing risks and improving precision [1].
@@
However, fully autonomous exploration inside underwater caves presents significant challenges: navigation is difficult without GPS, while poor visibility, backscattering, and silt disturbances hinder perception. Thus, uninformed exploration, as often done by AUVs in open-water ecological surveying and mapping tasks [31], is quite challenging. For autonomous navigation inside underwater caves [47, 66], it is important to be aware of the entry and exit directions as well as a semantic understanding of existing navigational markers, when available. More specifically, underwater caves explored by humans are augmented by a line (string), referred to as the caveline [14], which extends from the entrance through all the major sections of the cave. Other, navigational markers such as arrows and cookies [1] indicate the closest path to the exit (arrows’ directions) and the presence of landmarks and/or divers (cookies); real-life examples and demonstrations are provided in the the supplementary video. The caveline, together with the navigational markers, provides a 1D retraction of the 3D environment, indicating the orientation and distance from the entrance of the primary passages.
@@
In this paper, we demonstrate the system design and algorithmic integration of CavePI, a novel AUV specifically developed for navigating underwater caves using semantic guidance from cavelines and other navigational markers; see Fig. 1. Designed for single-person deployments, CavePI’s compact design and 4-degree-of-freedom (4-DOF) motion model enables safe traversal through narrow cave passages with minimal silt disturbance. It features a forward-facing camera for visual servoing and diver-robot interaction, and a downward-facing camera for caveline detection and tracking, effectively minimizing blind spots as low as $30^{\circ}$ below the robot, that converges at a distance of $38$ cm. A Ping sonar provides altitude data to maintain a safe distance from the ground plane. Additionally, two onboard lights enhance visibility in low-light conditions, complemented by on-device software visual filters [22] for improved perception. The computational framework of CavePI includes two single-board computers (SBCs): a Jetson Nano for perception, and a Raspberry Pi-5 for planning and control.
@@
For visual servoing, we integrate a caveline detector that performs real-time pixel-level segmentation of the caveline using the robot’s down-facing camera stream. While prior studies have proposed vision transformer (ViT)-based frameworks such as CL-ViT [66] and CaveSeg [1] for caveline detection, these approaches are computationally intensive and unsuitable for the resource-constrained SBCs. To address this, we adopt more efficient feature-extractor networks: MobileNetV3 [20] and ResNet101 [18] – combined with a lightweight DeeplabV3 [8] as the segmentation head. By leveraging GPU-accelerated computation on a Jetson Nano SBC, our system achieves a maximum inference rate of $18.2$ frames per second (FPS) while ensuring robust segmentation performance. The resulting segmentation map is post-processed to extract caveline contours, which are used to guide a proportional-integral-derivative (PID)-based Pure Pursuit controller [45] for precise heading adjustments for smooth tracking-by-detection [52]. Note that CavePI leverages the caveline position and directions for semantic guidance, while the ground plane is ranged by sonar. Other markers: arrows (exit direction) and cookies (divers’ presence) trigger momentary decisions such as taking u-turns, invoking logging threads, etc. Thus, in this paper, we mainly focused on demonstrating robust caveline following performances of CavePI.
@@
Following extensive in-house testing and refinement, CavePI is deployed in diverse real-world environments, including $1.5$-$6$ m deep spring waters and $12$-$30$ m deep natural underwater grottos and caves. These outdoor settings pose unique perception and navigation challenges, such as strong currents and turbidity in open waters, as well as noisy, low-light conditions within overhead environments. These deployments validate the robustness of CavePI, demonstrating its capability to perform long-term autonomous missions in complex underwater environments. Our long-term goal is to make it more generalizable for exploring any overhead structures (e.g., inside ship hulls, pipelines, dams) with no predefined markers.
[…]
@@ 2.1 Autonomous Underwater Cave Exploration
While SOTA AUVs are proficient in executing fully autonomous missions, their deployments have primarily been constrained to open water environments [48], where obstacles are minimal. Underwater caves, however, present unique and formidable navigation challenges, including overhead and surrounding obstacles, narrow passageways, and complex topologies. Early efforts to address these challenges include the work of Mallios et al. [30], where manually operated AUVs collected acoustic data for offline mapping. Similarly, Weidner et al. [63, 62] utilized stereo camera systems to generate high-resolution 3D reconstructions of cave walls, floors, and ceilings.
@@
Vision-based state estimation and navigation in underwater caves pose significant challenges due to factors such as lighting variability, backscattering effects, and image degradation [24]. To overcome these obstacles, Rahman et al. [42, 44] proposed a data fusion framework that integrates visual, acoustic, inertial, and water depth measurements, enabling robust trajectory estimation and sparse representations of underwater cave environments. Advances such as shadow-based mapping [43], contour extraction [33], and real-time stereo reconstruction [61] have further enhanced the density and fidelity of cave mapping techniques. More recently, visual learning-based approaches have shown promise for autonomous visual servoing within underwater caves [1, 66]. Other contemporary works with the SUNFISH AUV [47] have demonstrated effective navigation capabilities in these environments by employing Doppler Velocity Log (DVL)-based dead reckoning and sonar-based SLAM algorithms. However, it is not one-person portable, and requires considerable logistics for successful deployment.
@@ 2.2 Robot Navigation by Semantic Guidance
For navigating underwater caves safely, learning-based frameworks have been developed for detecting and following cavelines [66] and other navigation markers [1]. However, these computationally intensive models require further optimization to achieve real-time deployments [34]. Additionally, their onboard semantic segmentation performance and decision-making capabilities remain underexplored. We address this gap by conducting comprehensive real-world evaluations using our CavePI robotic platform, demonstrating its effectiveness for real-time semantic segmentation and navigation in complex underwater cave environments.
@@ 3.1 Sensor Bay: Acoustic-Optic Perception Subsystem
The CavePI platform includes visual and acoustic sensors: a front-facing fisheye camera, a downward-facing low-light camera, and a Ping2 active sonar. The fisheye camera, housed within a transparent dome at the head of the AUV, captures forward-facing visuals with a $160^{\circ}$ field-of-view (FOV) and outputs a video feed at $1920\times 1080$ resolution.This camera detects navigation markers and visual commands (QR codes) from divers, and will assist in obstacle avoidance in the next iteration. It is worth noting that the cylindrical enclosure is a $6^{\prime\prime}$ tube while the dome has a $4^{\prime\prime}$ diameter. A custom interface is built to connect the two; see section 5.1 for more details. The low-light camera, mounted inside the computational enclosure, captures downward-facing visuals with an $80^{\circ}\times 64^{\circ}$ FOV, also at the same resolution and frame rate. Additionally, a Ping2 sonar altimeter-echosounder from Blue Robotics™ is mounted on the underside of the robot; the sonar has a range of $100$ meters, a depth rating of $300$ meters, and a resolution of $0.5\%$ of the range, allowing it to detect obstacles directly beneath CavePI at an output frequency of $10$ Hz. These sensory components collectively provide robust environmental awareness for autonomous navigation in challenging underwater caves.
@@ 3.2 Computational Bay
The enclosure also contains the battery compartment, voltage regulators, electronic speed controllers (ESCs), and a Bar-30 pressure sensor. The battery compartment holds a $14.8$ V ($18$ Ah) battery pack, regulated to power internal components (e.g., cameras, computers) and external components (e.g., thrusters, sonar). The battery pack is capable of supplying sufficient power to sustain over 6 hours of operation of CavePI. Each thruster is controlled by an ESC, which drives the three-phase brushless motor using PWM signals from the Pixhawk. The Bar-30 sensor provides high-precision pressure readings with a resolution of $0.2$ mbar and an accuracy of $2$ mm, with a working depth of up to $300$ meters. This pressure data is processed to determine underwater depth, ensuring reliable and accurate interoceptive perception during operations.
@@ 3.3 Locomotion Bay: Middleware Integration
The CavePI AUV is designed for low-power operation and integrates a modular ROS2 framework to support application-specific perception, planning, and control methods. As depicted in Fig. 4, the detector node acquires visual data from the two cameras to identify the caveline for navigation. The mission planner node then integrates the caveline information with the estimated position data to generate subsequent waypoints for the mission. Finally, the autopilot controller node utilizes these waypoints, along with the detected caveline, positional data, and depth readings from the Bar-30 sensor, to generate precise actuation signals for the thrusters, enabling accurate movements and depth control. Additionally, CavePI can function as an ROV through an optional tether-based teleoperation module. This module transmits user commands from a joystick to the onboard Raspberry Pi-5, which processes the inputs and relays them to the thrusters for manual control.
Power footprint and synchronization. As shown in Table 1, CavePI offers an endurance of 6+ hours at maximum capacity. While it can operate much longer, the battery profile [11] recommends running it above $20$% power; the remaining capacity tends to deplete rapidly, which is a safety issue for recovery. While running in full capacity, the integrated system ensures an overall system throughput of $3.6$ Hz. Specifically, both the cameras run at $30$ Hz; the IMU and sonar frequencies are $100$ Hz and $10$ Hz, respectively. In our implementation, software synchronization is achieved via a joint ROS2 publisher that shares a clock across all sensory topics.
Table 1: The battery power consumption characteristics of CavePI (at maximum capacity) over time, showing 6+ hours of endurance.
Time (hours) 1 2 3 4 5 6 7
Battery % 85 73 61 50 39 29 19
@@ 3.4 CavePI Digital Twin
Beyond replicating caveline following experiments, we utilize the DT system for preliminary testing and fine-tuning of new control algorithms. It also enables the simulation of complex cave environments, such as narrow passages, dead ends, and sharp turns. Conducting repeated real-world experiments in such scenarios to improve the control system can be logistically demanding where the simulation offers an efficient alternative for extensive evaluation and fine-tuning.
@@ 4 CavePI Navigation Pipeline
The CavePI AUV is designed to autonomously navigate underwater by following a caveline and other navigation markers. However, a caveline appears a few pixels wide in the bottom-facing camera’s FOV and is significantly challenging to detect in noisy conditions with almost no ambient light. Moreover, the caveline is often obscured or blends into the background, requiring a robust visual learning pipeline for reliable operation.
@@ 4.1 Semantic Guidance: Tracking by Detection
The model selection process plays a crucial role in optimizing accuracy and efficiency on resource-constrained edge devices. While object detection and semantic segmentation are two established techniques in visual learning, they produce different interpretations of the target RoI. Object detection models generate bounding boxes around the target, which, given the caveline’s irregular shape and orientation, can encompass significant background regions, complicating the estimation of the heading angle. To address this, we opt for semantic segmentation, which provides pixel-level contours of the caveline as a more precise marker for tracking-by-detection.
@@
Model architecture and design choices. Onboard resource constraints of CavePI significantly influence the choice of model architecture. The Jetson Nano, with its shared memory architecture, allocates GPU memory dynamically from a limited $2$ GB pool, which must also accommodate the dual camera feeds and Nano-Pi control signal communication. Considering these limitations, we initially selected the MobileNetV3 (Large) backbone [20] for feature extraction. This lightweight architecture feeds extracted features into a DeepLabV3 [8] head to generate binary segmentation maps. While the MobileNetV3-DeepLabV3 combination is computationally efficient, requiring only about $314$ MB GPU memory for $11$ million parameters, it proved insufficient for accurately segmenting the thin caveline in turbid water conditions. To address this limitation, we adopted a heavier backbone, ResNet101 [18], while retaining the DeepLabV3 head. This model achieved slightly better performance, as highlighted in Table 2, albeit at the cost of reduced inference speed. Both architectures were initialized with pre-trained weights and fine-tuned on a custom dataset to enhance caveline detection accuracy in challenging environments.
Dataset and model training. For this study, we utilized the open-source CL-ViT dataset, which includes pixel-level annotations of cavelines [66]. It comprises $3,150$ RGB images collected from three distinct underwater cave systems. To enhance the dataset’s diversity and variance, we manually annotated an additional $150$ images from our laboratory testbed, resulting in a combined dataset of $3,300$ images. The dataset was divided into $3,200$ training and $100$ validation images. Both models were fine-tuned on this dataset over a maximum of $50$ epochs using the following training configurations: the Adam optimizer with an initial learning rate of $0.001$, a cosine learning rate scheduler with $T_{max}=20$, and a cross-entropy loss function. This training configuration was designed to optimize these models’ performances in diverse underwater environments.
Benchmark comparison. The fine-tuned models are evaluated on the open-source CL-Challenge dataset [66] that contains caveline annotations for $200$ noisy low-light images from different cave systems. Using the lighter MobileNetV3 backbone, our model achieves a mean intersection-over-union (mIoU) of $48.95\%$, whereas the considerably heavier ResNet101 model demonstrates only a marginal improvement with a mIoU of $50.48\%$. For reference, the highest mIoU score reported on this dataset is $58.3\%$ [66]. Since smooth real-time operation requires a minimum control loop of $3.6$ Hz, we opt for the MobileNetV3 backbone that exceeds this requirement with an inference speed of $18.2$ fps (frames per second).